OpenStudy (anonymous):

How to do Slope-intercept form? I have the straight up problem..just need to know how to graph it.

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

b=y-int m=slope

OpenStudy (anonymous):

plug in x and u get a point y

OpenStudy (anonymous):

The problem is y= x+5. how to do slope intercept?

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay..how did you know where to put the x coordinate? on the x line.